# SVInvNet: A Densely Connected Encoder-Decoder Architecture for Seismic Velocity Inversion

### v1

@@ Abstract
This study presents a deep learning-based approach to seismic velocity inversion problem, focusing on both noisy and noiseless training datasets of varying sizes. Our Seismic Velocity Inversion Network (SVInvNet) introduces a novel architecture that contains a multi-connection encoder-decoder structure enhanced with dense blocks. This design is specifically tuned to effectively process complex information, crucial for addressing the challenges of non-linear seismic velocity inversion. For training and testing, we created diverse seismic velocity models, including multi-layered, faulty, and salt dome categories. We also investigated how different kinds of ambient noise, both coherent and stochastic, and the size of the training dataset affect learning outcomes. SVInvNet is trained on datasets ranging from 750 to 6,000 samples and is tested using a large benchmark dataset of 12,000 samples. Despite its fewer parameters compared to the baseline, SVInvNet achieves superior performance with this dataset. The outcomes of the SVInvNet are additionally compared to those of the Full Waveform Inversion (FWI) method. The comparative analysis clearly reveals the effectiveness of the proposed model.
@@ I Introduction
The seismic reflection method is widely utilized to mapping subsurface structures, and it plays a crucial role in the exploration of hydrocarbon reservoirs and geothermal energy sources. In the surface-to-surface seismic data acquisition technique, an artificial source on the surface generates seismic energy that is transmitted underground. As this seismic wavefield propagates beneath the surface, it undergoes reflection, refraction, and diffraction at interfaces with varying seismic velocity values, causing it to return to the surface. The returning waves are captured by receivers placed on the surface and recorded as a time series. These recordings, known as seismic traces, are aligned next to one another and create seismic shot gather. This assembly of seismic shot gathers is instrumental in probing and understanding the properties beneath the Earth’s surface.
To extract depth-related information from time series seismic data, it’s essential to have a subsurface velocity distribution that can accurately predict the observed seismic data. The effectiveness of various seismic imaging techniques, including migration as detailed by Baysal et al. [1] and Berkhout[2] and also inversion methods as noted by Scales [3] and Schuster[4] depends on the precision of the velocity model. Although seismic data are crucial for developing the velocity model, their relationship is governed by the hyperbolic wave propagation equation, which is a nonlinear one. This complexity is compounded by the fact that different velocity models can yield identical seismic data, leading to an ill-posed inversion problem. This situation highlights the challenge of identifying a unique velocity model that aligns with the observed seismic data.
@@
In conjunction with ongoing investigations to address challenges associated with derivative-based methods, the effectiveness of innovative approaches is extensively examined. Significantly, within the realm of artificial intelligence advancements, deep neural networks stand out as a powerful tool for the mapping of seismic data to velocity model. The deep cascaded structure of these networks help learning complex non-linear functions from data. Through an iterative training process, the deep model adjusts the connections to capture the complex relationship between the seismic data and the velocity model. This approach allows the trained deep model to predict the velocity model accurately from seismic data, even if that data wasn’t part of the training set. This advancement reduces the reliance on initial velocity models, representing a major advancement in the efficiency and accuracy of seismic data interpretation.
@@
Due to the size and dimensional differences between seismic data and velocity model, the training process with deep models involves converting data from one domain to another. The encoder-decoder framework, which is good at deriving features from the input domain and predicting target in another domain, has demonstrated promising results in seismic data interpretation. In one of the earliest works, Wu et al.[11] used this architecture and designed an end-to-end convolutional neural network (CNN) called InversionNet. InversionNet was able to predict velocity models with four-layer horizontal interface and fault with certain success. Within the encoder segment of the InversionNet, the dimensions of the features undergo reduction to 1$\times$1, at the same time increasing the number of channels to 1024 progressively. Subsequently, the decoding segment tries to closely estimate the velocity model through the utilization of these 1$\times$1 features. The dimensions of the features in the penultimate layers are 112$\times$112, and the output layer constructs a 100$\times$100 velocity model via the interpolation method. The substantial incorporation of channel quantities, such as 512 and 1024, coupled with a feature size of 112$\times$112 in the latest layers, contributes to an elevation in the deep model’s parameter count to 44 million. Moreover, each layer is limited to utilize features exclusively from its immediate preceding layer.
Later, Li et al.[12] proposed SeisInvNet model using seismic data to construct velocity model. In SeisInvNet, which generally employs the encoder-decoder principle, each seismic trace is transformed into the features equivalent in size to the velocity model before entering the decoder section of the deep model. During this reshaping process, the initial step involves the conversion of input traces into embedding vectors. Subsequently, multiple fully connected layers operate to transfer this embedding vector into features matching the size of the velocity model. According to their approach, replacing the seismic trace with the embedding vector is beneficial in mitigating the spatial correspondence challenges between seismic data and the velocity model. The embedding vector is generated through the compilation of neighborhood information of the trace, extracted by a shallow CNN network. Additionally, the position of the receiver and source is integrated through the addition of a one-hot vector, and global context is derived via an encoder network.
@@
The improved version of SeisInvNet [13] were successful to predict multilayer velocity models including geologic structures such as fault and salt dome with inclined and undulating interfaces. In this study, receiver gathers are employed in addition to shot gathers as input data.
Our improved deep model, which we refer to as the Seismic Velocity Inversion Network (SVInvNet), features an end-to-end encoder-decoder design. Thanks to the multiple connections among its layers, the distance between input and ground truth is much shorter than in earlier designs. The topology of SVInvNet is based on learning the amplitude magnitude, the trajectory of the reflection/refraction/diffraction lined up near each others, and the arrival times of the seismic data, by CNN kernels, rather than increasing the number of channels. We utilize raw seismic traces within a simple and straightforward architecture, without applying complex transformations to the input.
@@
To effectively train deep networks for learning the complex nonlinear functions we’ve discussed, it’s crucial to have sufficient depth in the network. This depth is typically achieved by stacking multiple layers. However, increasing the depth introduces a challenge; vanishing gradient problem [14]. To tackle this problem, which is often encountered in deep CNN structures, we implement a variant of the DenseNet architecture, as introduced by Huang et al.[15]. Unlike traditional CNN architectures, DenseNet uniquely features multiple connections between each layer and its subsequent layers in a series of dense blocks, which effectively prevents the fading of derivatives and accelerates information flow. We propose a novel encoder-decoder type architecture where both the encoder and the decoder parts are designed using dense blocks. Furthermore, inspired from the UNet architecture, outputs from specific dense blocks in the encoder segment are connected to corresponding blocks in the decoder to support the information flow.
@@
In order to use in our research, we manually prepared a novel dataset with 18,000 labeled data, pairing seismic shot gathers with velocity models. To mimic real data more closely, we introduced both coherent and stochastic noise to the synthetic data, resulting in an additional 18,000 noisy labeled data. We experimented with five training datasets, each distinguished by variances in sample quantities from 750 to 6,000 samples. We assessed the impact of the data volume on our deep model’s learning capacity using a consistent test benchmark that we created for fair evaluations of the generalization capabilities of the models. The size of our test benchmark, comprising 12,000 pairs, is much larger in sample size than the training datasets.
@@
$\bullet$ We propose a novel end-to-end CNN-based encoder-decoder neural network architecture (SVInvNet) specifically designed for seismic velocity inversion task. SVInvNet is distinguished by its significantly reduced parameter count, thereby enhancing computational efficiency while maintaining high performance levels.
@@
$\bullet$ We establish a comprehensive benchmarking procedure for precise assessment and comparison of the trained deep learning models. Our approach uses a benchmark test set considerably larger than the training dataset, enhancing the evaluation of the model’s generalization capabilities in diverse scenarios.
$\bullet$ We provide a comprehensive analysis of the impact of varying training dataset sizes on the learning process for the seismic velocity inversion problem. To the best of our knowledge, this is the first time such an analysis has been presented in this domain.
@@ II Related Work
However, due to the limited computational facilities and insufficient development of the neural network method, this method has been ignored in seismic studies for many years. Recently, as advancements in computer technology and machine learning algorithms continue, especially in deep learning methodologies, we are witnessing a growing trend of deep network applications in the seismic exploration domain, consistent with developments observed in other scientific fields.
Lewis et al. [23] created a good initial velocity model for the FWI method to visualize the region under the salt dome using deep learning. For the training of the salt dome structure, they used real data (boreholes, well logs, and seismic data collected over many years) containing the salt dome structure in the hydrocarbon exploration region in the Gulf of Mexico. In this local study, post-stacking seismic gathers were used as the input.
@@ III Methodology and Implementation
Figure 1: Diagram illustrating (a) forward modeling and (b) inversion processes in seismic velocity studies utilizing deep learning-based models. The seismic propagation equation, $\mathcal{L}$, facilitates the computation of seismic data; however, no explicit equation is available for deriving velocity based on existing seismic data.
@@
Addressing this nonlinear ill-posed problem, researchers have explored a diverse range of methodologies. These span from global methods, which utilize Eq.(1) to identify an optimal $m_{v}$ by investigation in velocity model domain [27], to derivative-based approaches. The latter initiate with an initial model $\bar{m}_{v}$, subsequently compute the corresponding seismic data $\bar{D}$, and adjust the $\bar{m}_{v}$ parameters to minimize the local gradient of $L(D-\bar{D})$ [28], [5] (L is the loss function). More recent advancements have seen the emergence of deep learning-based techniques, providing innovative solutions to this long-standing problem [24], [11], [25], [12], [13].
The deep learning model can be mathematically represented as shown in Eq. (3):
@@
Here, $\mathcal{F}$ denotes the deep learning function responsible for transforming the seismic data, $D^{n}$, into a velocity model, $m^{n}_{v}$. $\mathcal{W}$ represents the set of learnable parameters or weights within the model. If N is the total number of training dataset, then each individual training sample can be denoted by n, such that n $\in$ N. The function $\mathcal{F}$ is composed of multiple neural network layers, each with varying depth and connection patterns, extending from the input to the output layer. The input data propagates through these layers to produce an output. This output is then compared to the ground truth or target and the distinction between them is quantified through the utilization of the loss function. The success rate of the deep model is influenced by both the quantity and comprehensiveness of the training dataset.
@@
In our research, we employed Convolutional Neural Networks (CNNs) based deep model to establish a clear link between the seismic data’s characteristics, such as direct arrivals, reflections, scattering, and potential refractions, and the underlying velocity model parameters (Fig. 1(b)) such as inter-layer velocity value, number of layers, layers interface shape, salt dome, and fault properties. To achieve this, we designed a neural network with a layered structure that is deep enough to maintain the integrity of the gradient flow during training. We introduced an encoder-decoder framework in our neural network, which allows us to train on complex models directly, bypassing the need for initial data manipulation or pre-processing. Furthermore, by integrating a Dense network structure, we were able to construct more robust neural network layers while ensuring that the gradient signal does not weaken significantly as it travels through the network. The details of this neural network design are elaborated in Section III-B.
@@ III-A1 Velocity Models
Two principles are considered in this process. First, the velocity value increases with depth, meaning that the lower layer always has a higher velocity than the upper one. Second, to partially maintain the geological sediment deposition logic between the layers, the shape of the interfaces above is influenced by those below. Consequently, the stochastically chosen interfaces merge with the ones below. The inter layer velocity value varies between [1500, 4000] m/s.
Algorithm 1 Layer Filling Algorithm
0:  Stochastic selection of parameters
@@
  $V_{i}\leftarrow$ Choose velocity value for the layer
  $d_{i}\leftarrow$ Choose depth of the layer
  $L_{i}\leftarrow$ Choose interface of the layer
@@ III-A2 Seismic Data
$\frac{\partial^{P}}{\partial t^{2}}-v^{2}(\frac{\partial^{2}P}{\partial x^{2}}+\frac{\partial^{2}P}{\partial z^{2}})=\mathcal{S}(t,x,z)$ (7)
@@
Figure 3: The velocity model and its associated synthetic shot gathers at the first (a), tenth (b), and twentieth (c) source locations.
@@ III-B Architecture of the Deep Model
In this study, we take the design of the InversionNet [11] deep model as the baseline. Similar to InversionNet, we employ one-dimensional (1D) CNN kernels to reduce the time axis of the input data to the number of receivers within the initial four layers. However, unlike InversionNet, we do not reduce the size of the features below 6$\times$6. This decision is based on the understanding that the amplitude of adjacent traces contains crucial information pertaining to the structure of the velocity model interfaces.
@@ III-B1 Dense Block Structure
Here, $L^{i}_{j}$ is the j-th layer of a dense block i, $x_{0}$ is input to the dense block and $y_{0}$ is the output of i-th dense block; $x^{i}_{j}$ is the output of the j-th layer. $c_{j}$ shows the channel size (i.e. number of feature maps generated), which determines the depth of the output volume, in each layer. In our architecture $c_{j}$= 64 and $y_{0}$ has 192 feature maps, and $L_{j}$ contain a 3$\times$3 convolution (Conv), batch normalization (BN) [33] layers, and rectified linear unit (ReLU)[34] activation function. For the sake of convenience, this combination, conv2d+BN+ReLU, will be referred to as the Con2D in our discussion from this on. In the transition layers between the dense blocks with identical feature sizes the channel size of the output is typically reduced (as shown with purple cubes in Fig. 6).

### v2

@@ Abstract
This study presents a deep learning-based approach to seismic velocity inversion problem, focusing on both noisy and noiseless training datasets of varying sizes. Our Seismic Velocity Inversion Network (SVInvNet) introduces a novel architecture that contains a multi-connection encoder-decoder structure enhanced with dense blocks. This design is specifically tuned to effectively process time series data, which is essential for addressing the challenges of non-linear seismic velocity inversion. For training and testing, we created diverse seismic velocity models, including multi-layered, faulty, and salt dome categories. We also investigated how different kinds of ambient noise, both coherent and stochastic, and the size of the training dataset affect learning outcomes. SVInvNet is trained on datasets ranging from 750 to 6,000 samples and is tested using a large benchmark dataset of 12,000 samples. Despite its fewer parameters compared to the baseline model, SVInvNet achieves superior performance with this dataset. The performance of SVInvNet was further evaluated using the OpenFWI dataset and Marmousi-derived velocity models. The comparative analysis clearly reveals the effectiveness of the proposed model.
@@ I Introduction
The seismic reflection method is widely utilized for mapping subsurface structures, and it plays a crucial role in the exploration of hydrocarbon reservoirs and geothermal energy sources. In the surface-to-surface seismic data acquisition technique, an artificial source on the surface generates seismic energy that is transmitted underground. As this seismic wavefield propagates beneath the surface, it undergoes reflection, refraction, and diffraction at interfaces with varying seismic velocity values, causing it to return to the surface. The returning waves are captured by receivers placed on the surface and recorded as a time series. These recordings, known as seismic traces, are aligned next to one another and create seismic shot gather. This assembly of seismic shot gathers is instrumental in probing and understanding the properties beneath the Earth’s surface.
To extract depth-related information from time series seismic data, it’s essential to have a subsurface velocity distribution that can accurately predict the observed seismic data. The effectiveness of various seismic imaging techniques, including migration as detailed by Baysal et al. [1] and Berkhout[2] and also inversion methods as noted by Scales [3] and Schuster[4] depends on the precision of the velocity model. Although seismic data are crucial for developing the velocity model, their relationship is governed by the hyperbolic wave propagation equation, which is nonlinear. This complexity is compounded by the fact that different velocity models can yield identical seismic data, leading to an ill-posed inversion problem. This situation highlights the challenge of identifying a unique velocity model that aligns with the observed seismic data.
@@
In conjunction with ongoing investigations to address challenges associated with derivative-based methods, the effectiveness of innovative approaches is extensively examined. Significantly, within the realm of artificial intelligence advancements, deep neural networks stand out as a powerful tool for the mapping of seismic data to velocity model. The deep cascaded structure of these networks help learning complex non-linear functions from data. Through an iterative training process, the deep model adjusts the connections to capture the complex relationship between the seismic data and the velocity model. This approach allows the trained deep model to predict the velocity model accurately from seismic data, even if that data wasn’t part of the training set [11], [12]. This advancement has the potential to reduce the reliance on initial seismic velocity models.
@@
Due to the size and dimensional differences between seismic data and velocity model, the training process with deep models involves converting data from one domain to another. The encoder-decoder framework, which is good at deriving features from the input domain and predicting target in another domain, has demonstrated promising results in seismic inversion. In one of the earliest works, Wu et al.[13] used this architecture and designed an end-to-end convolutional neural network (CNN) called InversionNet. InversionNet was able to predict velocity models with four-layer horizontal interface and fault with certain success. Within the encoder segment of the InversionNet, the size of the features undergo reduction to 1$\times$1, at the same time increasing the number of channels to 1024 progressively.
@@
Later, Li et al.[14] proposed SeisInvNet model using seismic data to construct velocity model. In SeisInvNet, which generally employs the encoder-decoder principle, each seismic trace is transformed into the features equivalent in size to the velocity model before entering the decoder section of the deep model. The improved version of SeisInvNet [15] were successful to predict multilayer velocity models including geologic structures such as fault and salt dome with inclined and undulating interfaces. In this study, receiver gathers are employed in addition to shot gathers as input data. In these architectures, each layer is limited to utilize features exclusively from its immediate preceding layer.
To effectively train deep networks for learning the complex nonlinear functions we’ve discussed, it’s crucial to have sufficient depth in the network. This depth is typically achieved by stacking multiple layers. In the training phase, the deep model attempts to estimate the velocity model by adjusting the weights in the layers. To facilitate this adjustment, it compares the result with the ground truth in each iteration and relays the difference through the layers to the first one using the backpropagation algorithm. Consequently, increasing the depth introduces a challenge; vanishing gradient problem [16]. To tackle this problem, which is often encountered in deep CNN structures, we implement a variant of the DenseNet architecture, as introduced by Huang et al.[17].
@@
Unlike traditional CNN architectures, DenseNet uniquely features multiple connections between each layer and its subsequent layers in a series of dense blocks, which effectively prevents the fading of derivatives and accelerates information flow between the layers and dense blocks. The feature size in each dense block remains constant, with the inputs to each layer formed by concatenating the outputs of all preceding layers. We propose a novel encoder-decoder type architecture where both the encoder and the decoder parts are designed using dense blocks.
Our improved deep model, which we refer to as the Seismic Velocity Inversion Network (SVInvNet) furthermore, inspired from the UNet architecture, outputs from specific dense blocks in the encoder segment are connected to corresponding blocks in the decoder to support the information flow. Thanks to the multiple connections among SVInvNet layers, the distance between input and ground truth is much shorter than in earlier designs. We utilize raw seismic traces within a simple and straightforward architecture, without applying complex transformations to the input.
@@
In order to use in our research, we manually prepared a novel dataset with 18,000 labeled data, pairing seismic shot gathers with velocity models. To mimic real data more closely, we introduced both coherent and stochastic noise to the synthetic data, resulting in an additional 18,000 noisy labeled data. We experimented with five training datasets, with varying sizes from 750 to 6,000. We assessed the impact of the data volume on our deep model’s learning capacity using a consistent test benchmark that we created for fair evaluations of the generalization capabilities of the models. The size of our test benchmark, comprising 12,000 pairs, is much larger in sample size than the training datasets. Additionally, two subsets of the synthetic dataset OpenFWI [18], which is openly available, are employed to assess the efficacy of the deep model we constructed. These subsets are known as CurveVel-B and CurveFault-A.
@@
$\bullet$ We propose a novel end-to-end CNN-based densely connected encoder-decoder neural network architecture (SVInvNet) specifically designed for seismic velocity inversion task. SVInvNet is distinguished by its significantly reduced parameter count, thereby enhancing computational efficiency while maintaining high performance levels.
@@
$\bullet$ We establish a comprehensive benchmarking procedure for precise assessment and comparison of the trained deep learning (DL) models. Our approach uses a benchmark test set considerably larger than the training dataset, enhancing the evaluation of the model’s generalization capabilities in diverse scenarios.
$\bullet$ We provide a comprehensive analysis of the impact of varying training dataset sizes on the learning process for the seismic velocity inversion problem.
@@ II Related Work
However, due to the limited computational facilities and insufficient development of the neural network method, this method has been ignored in seismic studies for many years. Recently, as advancements in computer technology and machine learning algorithms continue, especially in DL methodologies, we are witnessing a growing trend of deep network applications in the seismic exploration domain, consistent with developments observed in other scientific fields. In the domain of seismic inversion, you can find a timely overview of DL approaches up to 2020 in Adler et al. [26].
Lewis et al. [27] created a good initial velocity model for the FWI method to visualize the region under the salt dome using DL. For the training of the salt dome structure, they used real data (boreholes, well logs, and seismic data collected over many years) containing the salt dome structure in the hydrocarbon exploration region in the Gulf of Mexico. In this local study, post-stacking seismic gathers were used as the input.
@@
Research continues to advance in the pursuit of more reliable results. In the aforementioned studies, the estimation of the seismic velocity model has been approached as a regression problem. However, Simon et al. [31] employed both regression and classification methods to estimate the velocity model in the time domain in a comparative study. According to the results presented in their study, the classification method demonstrates a lower error rate in estimating velocity models compared to the regression method.
@@ III Methodology and Implementation
Figure 1: Diagram illustrating (a) forward modeling and (b) inversion processes in seismic velocity studies utilizing DL-based models. The seismic propagation equation, $\mathcal{L}$, facilitates the computation of seismic data; however, no explicit equation is available for deriving velocity based on existing seismic data.
@@
Addressing this nonlinear ill-posed problem, researchers have explored a diverse range of methodologies. These span from global methods, which utilize Eq.(1) to identify an optimal $m_{v}$ by investigation in velocity model domain [32], to derivative-based approaches. The latter initiate with an initial model $\bar{m}_{v}$, subsequently compute the corresponding seismic data $\bar{D}$, and adjust the $\bar{m}_{v}$ parameters to minimize the local gradient of $L(D-\bar{D})$ [33], [5] (L is the loss function). More recent advancements have seen the emergence of DL-based techniques, providing innovative solutions to this long-standing problem [28], [13], [29], [14], [15].
The DL model can be mathematically represented as shown in Eq. (3):
@@
Here, $\mathcal{F}$ denotes the DL function responsible for transforming the seismic data, $D^{n}$, into a velocity model, $m^{n}_{v}$. $\mathcal{W}$ represents the set of learnable parameters or weights within the model. If N is the total number of training dataset, then each individual training sample can be denoted by n, such that n $\in$ N. The function $\mathcal{F}$ is composed of multiple neural network layers, each with varying depth and connection patterns, extending from the input to the output layer. The input data propagates through these layers to produce an output. This output, in supervised method, is then compared to the ground truth or target and the distinction between them is quantified through the utilization of the loss function. The success rate of the deep model is influenced by both the quantity and comprehensiveness of the training dataset.
@@
In our research, we employed Convolutional Neural Networks (CNNs) based deep model to establish a clear link between the seismic data’s characteristics, such as direct arrivals, reflections, scattering, and potential refractions, and the underlying velocity model parameters (Fig. 1(b)) such as inter-layer velocity values, number of layers, layers interface shape, salt dome, and fault properties. To achieve this, we designed a neural network with a layered structure that is deep enough to maintain the integrity of the gradient flow during training. We introduced an encoder-decoder framework in our neural network, which allows us to train on complex models directly, bypassing the need for initial data manipulation or pre-processing. Furthermore, by integrating a Dense network structure, we were able to construct more robust neural network layers while ensuring that the gradient signal does not weaken significantly as it travels through the network. The details of this neural network design are elaborated in Section III-B.
@@ III-A1 Velocity Models
Two principles are considered in this process. First, the velocity value increases with depth, meaning that the lower layer always has a higher velocity than the upper one. Second, to partially maintain the geological sediment deposition logic between the layers, the shape of the interfaces above is influenced by those below. Hence, novel and stochastic interface configurations emerge, resulting in an expansion of the number of interfaces employed in the structure of velocity models beyond the initially designated 116 functions. Consequently, the stochastically chosen interfaces merge with the ones below. The inter layer velocity value varies between [1500, 4000] m/s.
Algorithm 1 Multi-Layer Filling Algorithm
0:  Stochastic selection of parameters, Total number of layers $N$
@@
  for $i=1$ to $N$ do
     $V_{i}\leftarrow$ Choose velocity value for the $i$-th layer
     $d_{i}\leftarrow$ Choose depth of the $i$-th layer
     $L_{i}\leftarrow$ Choose interface of the $i$-th layer
@@ III-A2 Seismic Data
$\frac{\partial^{2}P}{\partial t^{2}}-v^{2}(\frac{\partial^{2}P}{\partial x^{2}}+\frac{\partial^{2}P}{\partial z^{2}})=\mathcal{S}(t,x,z)$ (7)
@@
Figure 3: The velocity model and its associated synthetic shot gathers at the first, tenth, and twentieth source locations respectively, from left to right.
@@ III-B Architecture of the Deep Model
In this study, we take the design of the InversionNet [13] deep model as the baseline. Similar to InversionNet, we employ one-dimensional (1D) CNN kernels to reduce the time axis of the input data to the number of receivers within the initial four layers. However, unlike InversionNet, we do not reduce the size of the features below 6$\times$6. This decision is based on the understanding that the amplitude and the arrival times of adjacent traces contains crucial information pertaining to the structure of the velocity model interfaces. Various trials were conducted to determine the smallest size of the features in our deep model, with dimensions $4\times 4$, $5\times 5$, $6\times 6$, $7\times 7$, and $8\times 8$ being evaluated. Following the analysis of the results, dimension $6\times 6$ was chosen as it provided the best performance.
@@ III-B1 Dense Block Structure
Here, $L^{i}_{j}$ is the j-th layer of a dense block i, $x_{0}$ is input to the dense block and $y_{0}$ is the output of i-th dense block; $x^{i}_{j}$ is the output of the j-th layer. $c_{j}$ shows the channel size (i.e. number of feature maps generated), which determines the depth of the output volume, in each layer. In our architecture $c_{j}$= 64 and $y_{0}$ has 192 feature maps, and $L_{j}$ contain a 3$\times$3 convolution (Conv), batch normalization (BN) [38] layers, and rectified linear unit (ReLU)[39] activation function. For the sake of convenience, this combination, conv2d+BN+ReLU, will be referred to as the Conv2d in our discussion from this on. In the transition layers between the dense blocks with identical feature sizes the channel size of the output is typically reduced (as shown with purple cubes in Fig. 6). As evident from the design of the dense blocks, the extracted features are concatenated and used in the subsequent layers. Consequently, the input reaching these layers is richer in information. Additionally, the derivative information reaching the output part ($y_{0}$) of the dense block through backpropagation can be instantly distributed to all inner layers.
Figure 6: Proposed SVInvNet architecture for seismic velocity inversion. Best viewed in the electronic version with zoom.